OpenStudy (anonymous):

Find the area of a sector with a central angle of 32° and a radius of 8.5 millimeters. Round to the nearest tenth.

jimthompson5910 (jim_thompson5910):

What's the area of the circle with a radius of 8.5 mm

jimthompson5910 (jim_thompson5910):

Use A = pi*r^2

OpenStudy (anonymous):

Do I leave pi as is or do I actually incorporate it into the answer this time?

jimthompson5910 (jim_thompson5910):

well now it's saying to round to the nearest tenth, so I would go for 3.14 this time

jimthompson5910 (jim_thompson5910):

or you can use your calculator's pi button to get a longer version of pi

OpenStudy (anonymous):

it's fine I'm just suppose to round my answer off at the end, but the area for the circle would be 226.980

jimthompson5910 (jim_thompson5910):

good I'm getting about the same

jimthompson5910 (jim_thompson5910):

now if you wanted the complete circle, you'd be done, but you don't you only want a fraction of that area more precisely, you want 32/360 = 4/45 of this total area

jimthompson5910 (jim_thompson5910):

so multiply 4/45 by 226.980 and round to the nearest tenth to get your final answer

OpenStudy (anonymous):

awesome! I got 20.176 i.e. 20.2

jimthompson5910 (jim_thompson5910):

and I'm getting exactly the same, congrats

OpenStudy (anonymous):

Thank you so much n.n

jimthompson5910 (jim_thompson5910):

sure thing